# Fed-SB: A Silver Bullet for Extreme Communication Efficiency and Performance in (Private) Federated LoRA Fine-Tuning

### v1

@@ Appendix B Related Work
Differential Privacy (DP) and FL. A common limitation of standard FL frameworks is their susceptibility to privacy attacks, as clients publicly share model updates with a central server. To address this issue, DP is incorporated into FL methods to ensure the privacy of client updates. This work follows the approximate DP framework (Dwork, 2006; Dwork et al., 2014), which provides formal privacy guarantees for model updates. Privacy is enforced during training using the DP-SGD optimizer (Abadi et al., 2016), which applies gradient clipping and noise injection to protect individual contributions. Since DP is preserved under composition and post-processing (Dwork, 2006; Li et al., 2021), the final global model update also retains DP guarantees. Prior methods, such as Fed-IT and FedEx-LoRA, did not explicitly incorporate DP. This study extends these approaches to DP settings and benchmarks them alongside FFA-LoRA and the proposed method.
Appendix C Experiment Details
@@
We conduct experiments on a single NVIDIA A6000 GPU (48 GB) and report the average results from three independent runs. All non-private models are trained using the AdamW optimizer (Loshchilov and Hutter, 2019). In line with Ponkshe et al. (2024), we initialize the adapter matrices using just $1/{1000}$ of the respective training dataset size.
@@
Instruction Tuning. Table 6 presents the key hyperparameters and configurations for Mistral-7B, Gemma-2 9B, and Llama-3.2 3B. Our setup closely follows previous works (Hu et al., 2023; Ponkshe et al., 2024), ensuring consistency with established best practices. For the baseline experiments, we further set $\alpha=16$, consistent with prior literature (Singhal et al., 2025; Sun et al., 2024). We additionally perform a sweep over the learning rate for our experiments.
@@
(Federated) Private Fine-Tuning. Table 7 outlines the key hyperparameters and configurations for BERT-base in both centralized private and federated private settings. We train our models using the Opacus library (Yousefpour et al., 2021) with the DP-SGD optimizer (Abadi et al., 2016). Following standard DP practices, we set the privacy parameter as $\delta=\frac{1}{|\text{trainset}|}$. To ensure adherence to best practices, we adopt hyperparameter choices from prior works (Singhal et al., 2025; Hu et al., 2021). For baseline experiments, we additionally set $\alpha=16$, aligning with previous literature (Singhal et al., 2025; Sun et al., 2024). We additionally perform a sweep over the learning rate and maximum gradient norm in DP-SGD for our experiments.
@@
Dropout 00 00 00
# Clients $25$ $25$ $5$
Local Epochs $1$ $2$ $2$
Rounds $1$ $1$ $1$
@@
Epochs $3$
# Clients - $3$
Local Epochs - $6$
Rounds - $1$
Appendix D Dataset Details
CommonSense170K is a large-scale dataset that brings together eight benchmarks designed to assess various aspects of commonsense reasoning (Hu et al., 2023). Below is an overview of its constituent datasets:
@@ Appendix D Dataset Details
    PIQA (Bisk et al., 2020) evaluates physical commonsense by asking models to determine the most reasonable action in a given scenario.
@@
    ARC Easy (ARC-e) (Clark et al., 2018) consists of elementary-level science questions, serving as a fundamental test of a model’s reasoning abilities.
@@
    OBQA (Mihaylov et al., 2018) presents knowledge-intensive, open-book multiple-choice questions that require multi-step reasoning and retrieval.
@@
    HellaSwag (Zellers et al., 2019) tests contextual reasoning by asking models to predict the most plausible continuation of a passage from a set of candidates.
@@
    SIQA (Sap et al., 2019) examines social intelligence, requiring models to predict human actions and their social consequences.
@@
    ARC Challenge (ARC-c) (Clark et al., 2018) includes difficult multiple-choice science questions that demand deeper logical inference beyond statistical co-occurrence.
@@
    BoolQ (Clark et al., 2019) consists of naturally occurring yes/no questions, requiring models to infer relevant information from provided contexts.
@@
    WinoGrande (Sakaguchi et al., 2021) assesses commonsense knowledge through binary-choice sentence completion tasks that require resolving ambiguities.
The MetaMathQA dataset (Yu et al., 2024) constructs mathematical questions by reformulating them from different viewpoints while preserving their original knowledge content. We assess its performance using two well-established benchmarks: (1) GSM8K (Cobbe et al., 2021), a collection of grade-school-level math problems requiring step-by-step reasoning to reach a solution, and (2) MATH (Hendrycks et al., 2021), which consists of high-difficulty, competition-style problems designed to test advanced mathematical skills.
@@
Appendix E Additional Plots

### v2

@@ Appendix B Related Work
Differential Privacy (DP) and FL. A common limitation of standard FL frameworks is their susceptibility to privacy attacks, as clients publicly share model updates with a central server. To address this issue, DP is incorporated into FL methods to ensure the privacy of client updates. This work follows the approximate DP framework (17, 18), which provides formal privacy guarantees for model updates. Privacy is enforced during training using the DP-SGD optimizer (1), which applies gradient clipping and noise injection to protect individual contributions. Since DP is preserved under composition and post-processing (17, 31), the final global model update also retains DP guarantees. Prior methods, such as Fed-IT and FedEx-LoRA, did not explicitly incorporate DP. This study extends these approaches to DP settings and benchmarks them alongside FFA-LoRA and the proposed method.
Appendix C Initialization in Fed-SB
Fed-SB adopts the initialization strategy introduced in LoRA-SB to fix the adapter matrices $B$ and $A$. Proper initialization is crucial, since $B$ and $A$ remain frozen during training. For instance, if $B$ were initialized to zero (as in standard LoRA), the product $BRA$ would remain zero throughout, preventing any learning. In contrast, initializing $B$ and $A$ as orthonormal matrices ensures well-scaled gradients and allows Fed-SB to nearly match the performance of full fine-tuning.
To construct $B$ and $A$, we approximate the optimal update by averaging the first-step update across a small set of samples. A truncated SVD of this estimated update is then used to initialize the adapters. This requires only a small fraction of the training data (typically $0.1\%$), leading to negligible overhead in computation and time. Since the update is computed layerwise, memory usage during initialization never exceeds that of subsequent Fed-SB fine-tuning and remains below that of LoRA. Empirical analysis in LoRA-SB (39) shows that even $0.1\%$ of the samples is sufficient for stable initialization.
Appendix D Extensions to Rank-Heterogeneous Setting
In real-world federated deployments, client devices often operate under diverse computational budgets and memory constraints. This naturally leads to rank-heterogeneous settings, where different clients cannot train adapters of the same rank. Supporting such heterogeneity is important for practical adoption: while high-resource clients can benefit from richer low-rank subspaces, low-resource clients should still be able to participate meaningfully without being excluded from collaboration.
D.1 Rank-Heterogeneous Fed-SB
We extend Fed-SB to explicitly handle rank-heterogeneous clients while preserving its guarantees of exact aggregation. The key idea is to align all clients in a shared basis, chosen as the top $r_{\max}$ singular vectors of a reference weight matrix. Each client $i$ then selects a local rank budget $r_{i}\leq r_{\max}$ and optimizes within its most informative subspace:
$A_{i}=A[:,:r_{i}],\quad B_{i}=B[:r_{i},:],\quad R_{i}=R[:r_{i},:r_{i}].$
During aggregation, each client’s update $R_{i}$ is zero-padded (along rows and columns) to match the global dimension $r_{\max}\times r_{\max}$. This ensures that all updates are aligned in the same coordinate system and can be averaged exactly:
$R_{\text{agg}}=\frac{1}{c}\sum_{i=1}^{c}\text{pad}(R_{i}),\quad\Delta W=BR_{\text{agg}}A.$
In this formulation, low-rank clients contribute updates restricted to their subspaces, while high-rank clients provide richer information, and all updates combine seamlessly. Thus, Fed-SB can support heterogeneous client capabilities without loss of information, while maintaining exactness of aggregation.
D.2 Experiments
Table 7: Comparison of homogeneous and heterogeneous Fed-SB configurations for federated fine-tuning of Llama-3.2 3B on eight commonsense reasoning datasets.
Method BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Avg.
Homogeneous (all ranks = $120$) $64.86$ $81.66$ $74.87$ $81.67$ $75.22$ $86.03$ $70.56$ $72.25$ $75.89$
Heterogeneous (effective rank = $120$) $64.34$ $81.50$ $74.23$ $81.02$ $74.88$ $85.89$ $70.65$ $71.62$ $75.52$
Homogeneous (all ranks = $160$) $65.57$ $82.37$ $76.15$ $84.10$ $77.98$ $86.62$ $72.10$ $73.63$ $77.32$
Heterogeneous (effective rank = $160$) $64.83$ $82.05$ $76.43$ $83.92$ $77.53$ $85.96$ $71.90$ $72.98$ $76.95$
To assess the effectiveness of our federated rank-heterogeneous approach, we extend the commonsense reasoning experiments with Llama-3.2 3B to heterogeneous rank settings. For a fair comparison, we match the total rank budget of the homogeneous baselines ($120^{2}$ and $160^{2}$) by assigning client-specific ranks of $\{40,40,120,120,200\}$ and $\{60,60,180,200,220\}$, respectively. As shown in Table 7, Fed-SB achieves performance comparable to its homogeneous counterparts in both cases, demonstrating strong robustness to rank heterogeneity.
Appendix E Effect of Varying Rank on Fed-SB Performance
To further investigate the role of the rank parameter $r$, we conduct ablation studies of Fed-SB in both standard federated and privacy-preserving settings. In the non-private setting, we evaluate Mistral-7B and Gemma-2 9B fine-tuned on a subset of MetaMathQA across a wide range of rank values (r=32r=32240240), with results reported in Table 8. While selecting an optimal rank remains an open problem for all LoRA-based methods, our experiments show that intermediate values (r=120r=120200200) generally offer the best trade-off between performance and efficiency.
In the privacy-preserving setting, we evaluate centralized private Fed-SB using BERT-base fine-tuned on SNLI across ranks ranging from $16$ to $80$, with results presented in Table 9. Here, we observe that ranks in the range of 48488080 consistently achieve the strongest performance across different privacy budgets.
Overall, owing to Fed-SB’s lightweight design, we can scale to higher ranks when resources allow, yielding further performance improvements without incurring memory bottlenecks.
Table 8: Effect of varying Fed-SB rank ($r$) on federated fine-tuning performance of Mistral-7B and Gemma-2 9B, evaluated on GSM8K and MATH. Best results are in bold.
Rank Mistral-7B Gemma-2 9B
GSM8K ($\uparrow$) MATH ($\uparrow$) GSM8K ($\uparrow$) MATH ($\uparrow$)
$32$ $53.76$ $12.88$ $73.78$ $35.92$
$64$ $53.93$ $13.31$ $74.32$ $36.05$
$96$ $54.38$ $13.56$ $74.66$ $36.23$
$120$ $54.44$ $\mathbf{14.06}$ $74.75$ $36.36$
$160$ $54.81$ $13.74$ $76.88$ $36.94$
$200$ $56.18$ $13.76$ $77.03$ $\mathbf{37.56}$
$240$ $\mathbf{56.32}$ $13.74$ $\mathbf{77.14}$ $37.34$
Table 9: Effect of varying Fed-SB rank ($r$) on centralized private fine-tuning performance of BERT-base, evaluated on SNLI, under various privacy budgets ($\epsilon$). A smaller $\epsilon$ indicates a stricter privacy budget. Best results are in bold.
Rank Accuracy ($\uparrow$)
$16$ $73.26$ $74.21$ $73.68$ $76.23$ $75.80$
$24$ $73.65$ $74.78$ $73.92$ $76.88$ $76.02$
$32$ $73.99$ $75.09$ $74.45$ $77.01$ $76.24$
$48$ $\mathbf{75.98}$ $75.70$ $76.58$ $76.77$ $77.96$
$64$ $75.81$ $\mathbf{77.07}$ $\mathbf{77.59}$ $78.75$ $78.08$
$80$ $75.93$ $76.87$ $77.35$ $\mathbf{78.81}$ $\mathbf{78.23}$
Appendix F Memory and Training Time Details
Memory. As discussed in Section 4.3, our method reduces training memory requirements compared to existing approaches, primarily due to a significantly smaller number of trainable parameters. We benchmark the peak per-client training memory for all models and configurations used in our study in Table 10. Notably, these results reflect the worst-case setting for Fed-SB, with the highest rank ($r=200$) used in our experiments.
Table 10: Peak per-client training memory (in GB) for different methods across the various models used in this work. Fed-SB consistently exhibits lower memory usage across all model configurations.
Method Rank Peak Memory (GB)
Mistral-7B Gemma-2 9B Llama-3.2 3B
FedIT $32$ $15.92$ $19.99$ $7.71$
FFA-LoRA $32$ $15.51$ $19.44$ $7.46$
FedEx-LoRA $32$ $15.92$ $19.99$ $7.71$
FLoRA $32$ $15.92$ $19.99$ $7.71$
Fed-SB $200$ $15.18$ $19.03$ $7.30$
Training Time. Fed-SB introduces a negligible training time overhead compared to other methods, primarily due to its lightweight initialization process. To quantify this, we measure the additional training time introduced by Fed-SB relative to the average per-epoch training time per client in baseline methods. These measurements are conducted across the various experimental settings described in our paper. As shown in Table 11, the overhead remains consistently minimal, approximately $2\%$, across multiple model configurations.
Table 11: Training time overhead introduced by Fed-SB ($r=200$) relative to the average per-epoch training time per client in baseline methods. The overhead is minimal ($\approx 2\%$) across different model configurations.
Model Fed-SB Overhead (mm:ss) Avg. Epoch Time / Client (mm:ss)
Mistral-7B 00:13 09:22
Gemma-2 9B 00:16 12:43
Llama-3.2 3B 01:43 62:54
Appendix G Experiment Details
@@
We conduct experiments on a single NVIDIA A6000 GPU (48 GB) and report the average results from three independent runs. All non-private models are trained using the AdamW optimizer (33). To optimize memory efficiency, all base models (except BERT) are loaded in torch.bfloat16. In line with LoRA-SB (39), we initialize the adapter matrices using just 1/10001/{1000} ($0.1\%$) of the respective training dataset size.
@@
Instruction Tuning. Table 12 presents the key hyperparameters and configurations for Mistral-7B, Gemma-2 9B, and Llama-3.2 3B. Our setup closely follows previous works (22, 39), ensuring consistency with established best practices. For the baseline experiments, we further set $\alpha=16$, consistent with prior literature (43, 44). We additionally perform a sweep over the learning rate for our experiments.
@@
(Federated) Private Fine-Tuning. Table 13 outlines the key hyperparameters and configurations for BERT-base in both centralized private and federated private settings. We train our models using the Opacus library (53) with the DP-SGD optimizer (1). Following standard DP practices, we set the privacy parameter as $\delta=\frac{1}{|\text{trainset}|}$. To ensure adherence to best practices, we adopt hyperparameter choices from prior works (43, 21). For baseline experiments, we additionally set $\alpha=16$, aligning with previous literature (43, 44). We additionally perform a sweep over the learning rate and maximum gradient norm in DP-SGD for our experiments.
@@
Dropout 0 0 0
# Clients $25$ $25$ 55
Local Epochs 11 22 22
Rounds 11 11 11
@@
Epochs 33 -
# Clients - 33
Local Epochs - 66
Rounds - 11
Appendix H Dataset Details
CommonSense170K is a large-scale dataset that brings together eight benchmarks designed to assess various aspects of commonsense reasoning (22). Below is an overview of its constituent datasets:
@@ Appendix H Dataset Details
    PIQA (6) evaluates physical commonsense by asking models to determine the most reasonable action in a given scenario.
@@
    ARC Easy (ARC-e) (12) consists of elementary-level science questions, serving as a fundamental test of a model’s reasoning abilities.
@@
    OBQA (37) presents knowledge-intensive, open-book multiple-choice questions that require multi-step reasoning and retrieval.
@@
    HellaSwag (56) tests contextual reasoning by asking models to predict the most plausible continuation of a passage from a set of candidates.
@@
    SIQA (42) examines social intelligence, requiring models to predict human actions and their social consequences.
@@
    ARC Challenge (ARC-c) (12) includes difficult multiple-choice science questions that demand deeper logical inference beyond statistical co-occurrence.
@@
    BoolQ (11) consists of naturally occurring yes/no questions, requiring models to infer relevant information from provided contexts.
@@
    WinoGrande (41) assesses commonsense knowledge through binary-choice sentence completion tasks that require resolving ambiguities.
The MetaMathQA dataset (55) constructs mathematical questions by reformulating them from different viewpoints while preserving their original knowledge content. We assess its performance using two well-established benchmarks: (1) GSM8K (13), a collection of grade-school-level math problems requiring step-by-step reasoning to reach a solution, and (2) MATH (20), which consists of high-difficulty, competition-style problems designed to test advanced mathematical skills.
@@
Appendix I Additional Plots
@@
Appendix J Use of Large Language Models
Our use of LLMs is limited to minor writing assistance, for example, correcting grammar and clarifying sentences.